# Bayesian and Multi-Objective Decision Support for Real-Time Cyber-Physical Incident Mitigation

### v1

@@ 3.2 Exposure Calculations
In the running example, EPSS scores corresponding to the CVE identifiers of vulnerability nodes are retrieved from the FIRST EPSS database and assigned as exposure probabilities. There is however one vulnerability node, V16, that does not have a CVE identifier. In this case, we approximate the CVSS vector for V16, which denotes a dynamic load-changing attack on the power grid, based on information from the literature [9]:
@@
where: AV(Attack Vector):N(etwork) indicates remote exploitation over a network; AC(Attack Complexity):L(ow) reflects that no specialised conditions are required; PR(Privileges Required):N(one) and UI(User Interaction):N(one) denote that the attack requires neither privileges nor user interaction. S(Scope):C(hanged) implies that the impact extends beyond a single system, affecting the broader power grid. C(Confidentility):N(one) indicates no confidentiality loss;
I(Integrity):L(ow) reflects partial integrity compromise through unauthorised control; and A(Availabaility):H(igh) denotes a high impact on system availability, potentially leading to widespread disruption.
Using Equation 2, the fallback exposure probability estimate for V16 is calculated as:
$P(E)_{\text{V16, CVSS}}=0.85\times 0.77\times 0.85\times 0.85=0.4729$ (12)
Next, we apply our Bayesian confidence calibration technique to the fallback estimate $P(E)_{\text{V16, CVSS}}$ to improve its realism and robustness. First, a standard deviation of 0.2 is assigned to reflect the low confidence level of the CVSS-proxy estimation, yielding a likelihood variance of $\sigma_{\text{CVSS}}^{2}=0.04$. For the prior distribution, a mean of $\mu_{0}=0.5$ and a variance of $\sigma_{0}^{2}=0.0025$ (corresponding to a standard deviation of 0.05) are assigned to reflect a neutral initial belief in the distribution. Substituting these values into Equations 7 and 8 yields the following:
@@
Additionally, an AF modifier of 1 is applied to all vulnerability node exposure probabilities using Equation 9 to reflect an idealised modelling assumption in which the adversary possesses unrestricted access and capability, and the target system exhibits minimal defensive barriers.
@@ 3.3 Impact Probability Calculations
In the case of vulnerability nodes, we use the following formulation to compute the probability of impact [15]:
@@
In the running example, the impact probabilities of vulnerability nodes are calculated using Equation 15 with corresponding confidentiality, integrity, and availability impact metrics drawn from the CVSS vectors. The impact probabilities of asset and hazard nodes are calculated using Equation 16 with the corresponding values derived from the BN graph.
@@ 3.4 Posterior Probability Calculations
Once the exposures and impact probabilities for vulnerability, asset, and hazard nodes are calculated, the posterior likelihood of successful attack and the probability of severe impact to the CPS – given the evidence of a vulnerability node being exploited, for example – can be derived using conditional probability formulations and Variable Elimination (VE) techniques [18].
@@
For the running example, the joint conditional probabilities for exposure and impact associated with the attacker’s goal: H8_Power_Outage are shown in Table 1. These values are derived from Bayesian inference computations based on estimates, posterior parameters, and updated evidence from the risk model.
@@
The results indicate that the posterior likelihood of a successful attack (exposure) is $P(E)=0.2555$, while the posterior probability of severe impact on the CPS is $P(I)=0.9153$. Multiplying these yields a composite risk score of $R=P(E)\times P(I)=23.39\%$, highlighting the potential for high-impact outcomes even when the estimated probability of attack success is relatively moderate.
@@ 3.5 Multi-Objective Optimisation
Our framework employs multi-objective optimisation to assess and weigh trade-offs between countermeasure portfolios, failure probabilities, and impact ratings to inform consequence-driven and risk-informed decision support in CPS.
@@
The first author’s experience suggests that asset failure probabilities can be influenced by mitigation-induced risks. For instance, the application of faulty or unstable security patches to a CPS asset may introduce unintended consequences, including asset failure. This complexity is exacerbated when an asset is linked to multiple vulnerabilities, each associated with distinct mitigation-related risks. In this framework, each countermeasure is associated with a risk adjustment function that governs its net contribution to asset-level failure. Accordingly, the failure probability of a CPS asset is defined as a bounded, risk-adjusted function of the mitigations applied across its associated vulnerabilities:
@@
Another important element in the framework is the Impact
Rating attribute, embedded in asset and hazard nodes. This rating captures the multifaceted consequences of successful attacks across safety, operational, financial, and informational domains. It is computed using an adaptation of the formulation proposed by Amro et al. [3]:
$\texttt{Impact Rating}=\sum_{j\in\{S,F,I,O,C\}}\left(\text{Factor}_{j}\cdot\text{Criticality}_{j}\right)$ (18)
@@
  • F: financial impact — quantifies monetary losses due to CPS equipment damage, service downtime, or disrupted business operations.
  • I: informational impact — relates to the compromise, loss, or exposure of sensitive data and system configurations.
@@
  • C: staging impact — indicates the ability to enable further attacks or facilitate persistence, privilege escalation, or lateral movement.
@@
Each portfolio’s mitigation values are encoded into the BN model and used to infer the posterior probabilities of both attack success and severe impact on the CPS, via Bayesian inference algorithms. These probabilistic outcomes are treated as parameters within a multi-objective search space. Pareto-optimal portfolios are then identified using the Optuna library [1], balancing the competing objectives of minimising the likelihood of attack success, and at the same time, minimising the probability of severe impact and maximising CPS availability. The resulting Pareto fronts form the foundation of our proposed decision support mechanism.
@@
In the running example, Equation 17 is used to calculate the risk-adjusted failure probabilities of CPS assets such as the WiNet-S communication dongle, which are connected to multiple vulnerability nodes, each associated with distinct mitigation-related risks.
Equation 18 is used to calculate impact ratings. As shown in Table 2, the impact rating of the solar PV inverter is calculated to be 0.3864. This figure arises from assigning a maximum safety factor of 1.0, based on the overarching principle that safety must never be compromised, although the associated criticality is 0, as no injuries are anticipated in the event of failure. Financial impact carries a weight of 0.25 and a criticality of 0.25, signifying localised equipment damage. The informational dimension, also weighted at 0.25, presents a higher criticality of 0.75 owing to the component’s role in managing system configurations that could substantially affect overall CPS security. Operationally, the component is heavily weighted (0.75) with a matching criticality of 0.75, as it directly influences CPS operations – its failure potentially leading to cascading outages or grid instability. Finally, the staging impact holds a moderate weight of 0.5 and a criticality of 0.5, denoting the component’s potential to facilitate attacker persistence or lateral movement.
@@
Finally, a multi-optimisation study is performed to identify Pareto-optimal countermeasure portfolios. The scatter plot (Figure 3) reveals a consistent curved slope across the likelihood and impact axes, with data points forming a relatively smooth trajectory that appears to ascend with increasing impact levels. This suggests a stable correlation between these dimensions, wherein higher-impact and higher-availability scenarios tend to co-occur. The continuity of the slope implies underlying systemic regularities, potentially driven by latent causal BN-represented structures within the CPS environment. The scatter plot also displays a plateau between approximately 0.25 and 0.30 along the Impact axis. This observation indicates that a subset of countermeasure portfolios successfully reduce the likelihood of severe impact, yet offer limited gains in CPS availability.
@@ 3.6 Frequency-Based Heuristic Analysis
Lastly, a frequency-based heuristic analysis of Pareto-optimal portfolios is conducted across multiple optimisation runs to rank and identify the vulnerability nodes that most consistently and effectively contributed to high-performing countermeasure configurations. These nodes are strong candidates for prioritisation in CPS mitigation strategies, particularly under time and resource constraints during incident response.
@@
As shown in Figure 4, vulnerabilities V1 through V4 demonstrate consistently high mitigation effectiveness across numerous portfolio instances. This finding is corroborated by the underlying Bayesian Network structure (Figure 2), which reveals that these vulnerabilities constitute critical entry points, facilitating downstream attack propagation throughout the CPS.
@@
  • Optimisation results reveal systemic regularity, suggesting latent causal structures within the CPS BN structure.
@@
  • A subset of countermeasure portfolios successfully reduce the likelihood of severe impact, yet offer limited gains in CPS availability.
@@
  • Vulnerabilities V1V4 demonstrated consistently high mitigation effectiveness, reflecting their positions as critical entry points in the CPS BN structure.
@@
Taken together, these findings point to several practical implications for CPS practitioners. The observed systemic regularity invites further investigation into causal structures, potentially enabling more targeted interventions. Furthermore, the limited availability gains observed suggest the need for integrating complementary strategies such as anomaly detection, automated containment, or real-time system reconfiguration to strengthen operational resilience. The consistent effectiveness of vulnerabilities V1V4 reinforces their prioritisation as part of any initial mitigation sweep, especially in scenarios involving resource-constrained decision-making.
4 Evaluation and Results
@@
This section evaluates the practical utility and applicability of our proposed framework by applying the methodology to another two representative CPS scenarios beyond the initial case study. The corresponding results are presented and analysed to highlight key insights, limitations, and implications for real-world deployment.
4.1 Ukrainian Power Grid Attack (BlackEnergy)
@@
Our proposed framework is applied to enable real-time, adaptive mitigation of the 2015 Ukrainian power grid attack, which notoriously leveraged the BlackEnergy malware to attack a CPS [34]. Figure 5 presents the corresponding Bayesian Network representation of the attack tree [27] created with the GeNIe Academic tool [11]. Based on this structure, the BN model is constructed using available information to define the asset, vulnerability, and hazard nodes, and encoded in AutomationML [4].
Figure 5: Bayesian Network representation of the BlackEnergy cyber attack, adapted from the attack tree model proposed by Kumar et al. [27]. Nodes shaded in yellow denote CPS assets; blue nodes represent vulnerabilities; and purple nodes correspond to hazards. The grey-shaded node indicates attack feasibility, whilst the red-shaded node reflects the attacker’s objective, which is to disrupt electricity supply.
@@
Next, we calculate the exposure and impact probability values using the methodology outlined in Section 3. Two of the vulnerability nodes are listed in the National Vulnerability Database (NVD) [31] and associated with CVE identifiers, whereas the remaining vulnerability nodes are not. As a fallback, the exposure probabilities of non-CVE vulnerabilities must be estimated and calibrated. Taking the example of one such vulnerability node V3, which represents a malware downloader, the CVSS-proxy vector is approximated as:
@@
This CVSS vector characterises a local attack scenario within the CPS: AV:L indicates the attacker requires local access to the target asset; AC:L denotes low complexity with no specialised preconditions; PR:N and UI:N indicate that exploitation requires no privileges or user interaction. S:U shows the impact is limited to the same system. The C:H, I:H, and A:H designations reflect high levels of confidentiality, integrity, and availability loss, respectively, implying complete system compromise, including access to sensitive data, arbitrary control, and potential operational failure.
Based on the CVSS vector, the exposure probability estimate is derived as:
$P(E)_{\text{vuln, CVSS}}=0.306$ (19)
@@
Table 3 presents the conditional probabilities for exposure and impact associated with the attacker’s goal: H5_Disable_Electrical_Supply, derived using Bayesian inference via a Variable Elimination algorithm. The results show a posterior likelihood of successful attack of $P(E)=0.875$, and a corresponding posterior probability of severe impact on the CPS of $P(I)=0.2393$. The composite risk score is computed as $R=P(E)\times P(I)=20.94\%$ in the BlackEnergy attack.
@@
Next, we performed optimisation studies to search for Pareto-optimal countermeasure portfolios to support decision-making in mitigating the BlackEnergy attack. Figure 6 reveals an increasingly linear relationship between impact probability and CPS availability at higher impact probability levels. Additionally, the Pareto fronts appear clustered near a similar value along the attack likelihood axis, suggesting limited variation in attack likelihood across trials.
@@
Inspection of the top-performing countermeasure portfolio (Table 4) reveal near-maximal intervention against all vulnerabilities, with all mitigation parameters close to 1.0. Despite this, the probability of severe impact remained moderate. This suggests that in the case of the BlackEnergy attack, aggressive mitigation improves system availability, but does not fully suppress the residual risk of severe impact on the CPS.
@@
In summary, the key findings from the framework’s application to the Ukranian power grid attack are:
  • Increasingly linear relationship between impact probability and CPS availability at higher impact probability levels.
@@
  • Aggressive mitigation improves availability, but does not fully suppress the residual risk of severe impact on the CPS.
@@
These insights equip CPS practitioners with the rationale to adopt a multi-pronged mitigation strategy tailored to the dynamics of BlackEnergy-like threats. Firstly, CPS practitioners should invest in early-stage interventions to reduce exposure before the system enters high-risk zones. Secondly, the clustering of Pareto fronts around similar attack likelihood values calls for prioritising interventions based on impact severity and system availability, rather than likelihood alone. Even when near-total mitigation coverage is pursued aggressively, residual risks may persist. This warrants the need to complement countermeasure portfolios with real-time anomaly detection, automated containment mechanisms, or fail-safe protocols to arrest cascading failures. Finally, CPS practitioners should adopt mitigation strategies designed to target multiple vulnerabilities in parallel to enhance CPS resilience.

### v2

@@ 3.2 Exposure Probability Calculations
In our running example of a solar PV inverter, EPSS scores and KEV catalog entries for CVE identifiers of vulnerability nodes are retrieved from the FIRST EPSS and CISA databases respectively. LEV scores are computed per Equation 1 using EPSS scores over a 30-day window ($w=30$). For vulnerability node V8 (CVE-2024-50687: iSolarCloud API allows unauthorised access to Sungrow device data and parameters), these metrics are presented in Table 1. Equation 2 then yields a final exposure probability of $P(E)_{V8,CVE}=0.0342$.
Table 1: Vulnerability metrics for CVE-2024-50687 (retrieved 2026-02-15). Despite a high CVSS score (9.1), the LEV-derived exposure probability remains low (0.0342), indicating that although the vulnerability has a high theoretical severity score, it is unlikely to be exploited in real-world scenarios.
Scoring System Score
CVSS v3.1 9.1
EPSS (Latest Score) 0.0012
LEV 0.0342 (3.42%)
KEV 0 (Not listed in CISA catalogue)
$P(E)_{\text{V8, CVE}}$ 0.0342
For the single vulnerability node V16, which does not have a CVE identifier and represents a dynamic load-changing attack on the power grid, we approximate the CVSS-proxy vector based on information from the literature [10].
@@
This CVSS vector characterises a network-accessible attack within the critical infrastructure: AV:N indicates remote exploitation over a network; AC:L denotes low complexity requiring no specialised preconditions; PR:N and UI:N show that no privileges or user interaction are needed; S:C reflects changed scope impacting beyond the target system; C:N indicates no confidentiality loss; I:L shows low integrity compromise; while A:H signifies high availability impact. Using Equation 3 with the corresponding CVSS metric values [16], the fallback exposure probability estimate for V16 is calculated as:
$P(E)_{V16,non-CVE}=0.85\times 0.77\times 0.85\times 0.85=0.4729$ (10)
Next, we apply our Bayesian confidence calibration technique to the fallback exposure probability estimate $P(E)_{V16,non-CVE}$ to improve its realism and robustness. First, a standard deviation of 0.2 is assigned to reflect the low confidence level of the CVSS-proxy estimation, yielding a likelihood variance of $\sigma_{\text{CVSS}}^{2}=0.04$. For the prior distribution, a mean of $\mu_{0}=0.5$ and a variance of $\sigma_{0}^{2}=0.0025$ (corresponding to a standard deviation of 0.05) are assigned to reflect a neutral initial belief in the distribution. Substituting these values into Equations 5 and 6 yields the following:
@@
Additionally, an AF modifier of $\phi=1.0$ is applied to all vulnerability node exposure probabilities using Equation 7 to reflect an idealised modelling assumption in which the adversary possesses unrestricted access and capability, and the target system exhibits minimal defensive barriers.
@@ 3.3 Impact Probability Calculations
In the case of vulnerability nodes (both CVE- and non-CVE-linked), we use the following formulation, drawing on the corresponding CVSS metric values [16], to compute the probability of impact:
@@
In our running example, the impact probabilities of vulnerability nodes are calculated using Equation 13, drawing on the confidentiality, integrity, and availability metric values derived from their respective CVSS vectors. For instance, the vulnerability node V8 (CVE-2024-50687) has a CVSS vector: CVSS:3.1/AV:N/AC:L/PR:N/UI:N/S:U/C:H/I:H/A:N. Here, C:H indicates high confidentiality loss (metric value: 0.56); I:H shows high integrity compromise (metric value: 0.56); and A:N signifies no availability impact (metric value: 0). Accordingly, the impact probability for this vulnerability node is calculated as follows:
$P(Impact)_{V8}=1-[(1-0.56)\times(1-0.56)\times(1-0)]=0.8064$ (15)
Within the solar PV inverter BN presented in Figure 2, which comprises 30 nodes in total, the impact probabilities of hazard and asset nodes are derived from the underlying interdependencies among them. For illustration, successful exploitation of an asset node app, specifically the Sungrow iSolarCloud mobile/web application, directly influences three vulnerability nodes: V1, V2, and V3. Using Equation 14, this asset node’s impact probability is calculated as follows:
$P(Impact)_{\textit{app}}=\frac{3}{30}=0.1$ (16)
@@ 3.4 Posterior Probability Calculations
Once the exposure and impact probabilities for vulnerability, asset and hazard nodes are calculated, the posterior likelihood of successful attack and the probability of severe impact to the CPS, given the evidence of a vulnerability node being exploited, for example, can be derived using conditional probability tables and variable elimination [49].
@@
For the running example, the joint conditional probabilities for exposure and impact associated with the attacker’s goal: H8_Power_Outage are shown in Table 2. These values are derived from Bayesian inference computations based on estimates, posterior parameters and updated evidence from the risk model.
@@
The results indicate that the posterior likelihood of a successful attack (exposure) is $P(E)=0.2555$, while the posterior probability of severe impact on the CI is $P(I)=0.9153$. Multiplying these yields a composite risk score of $R=P(E)\times P(I)=23.39\%$, highlighting the potential for high-impact outcomes even when the estimated probability of attack success is relatively moderate.
@@ 3.5 Multi-Objective Optimisation
Our framework employs multi-objective optimisation to assess and weigh trade-offs between countermeasure portfolios, failure probabilities and impact ratings to inform consequence-driven and risk-informed decision support in CPS.
@@
The first author’s professional experience suggests that asset failure probabilities can be influenced by mitigation-induced risks. For instance, the application of faulty or unstable security patches to a CPS asset may introduce unintended consequences, including asset failure. This complexity is exacerbated when an asset is linked to multiple vulnerabilities, each associated with distinct mitigation-related risks. In our framework, each countermeasure is associated with a risk adjustment function that governs its net contribution to asset-level failure. Accordingly, the failure probability of a CPS asset is defined as a bounded, risk-adjusted function of the mitigations applied across its associated vulnerabilities:
@@
Another important element in the framework is the Impact Rating attribute, embedded in asset and hazard nodes. This rating captures the multifaceted consequences of successful attacks across safety, operational, financial and informational domains. It is computed by adapting the formulation proposed by Amro et al. [3]:
$\text{Impact Rating}=\sum_{j\in\{S,F,I,O,C\}}\left(\text{Factor}_{j}\cdot\text{Criticality}_{j}\right)$ (18)
@@
  • F: financial impact — quantifies monetary losses due to CPS equipment damage, service downtime or disrupted business operations.
  • I: informational impact — relates to the compromise, loss or exposure of sensitive data and system configurations.
@@
  • C: staging impact — indicates the ability to enable further attacks or facilitate persistence, privilege escalation or lateral movement.
@@
Each portfolio’s mitigation values are encoded into the BN model and used to infer the posterior probabilities of both attack success and severe impact on the CI, via Bayesian inference algorithms. These probabilistic outcomes are treated as parameters within a multi-objective search space. Pareto-optimal portfolios are then identified using the Optuna library [1], balancing the competing objectives of minimising the likelihood of attack success, and at the same time, minimising the probability of severe impact and maximising system availability. The resulting Pareto fronts form the foundation of our proposed decision support mechanism.
@@
In our running example, Equation 17 is used to calculate the risk-adjusted failure probabilities of CPS assets such as the WiNet-S communication dongle, which are connected to multiple vulnerability nodes, each associated with distinct mitigation-related risks.
Equation 18 is used to calculate impact ratings. As shown in Table 3, the impact rating of the solar PV inverter is calculated to be 0.3864. This figure arises from assigning a maximum safety factor of 1.0, based on the overarching principle that safety must never be compromised, although the associated criticality is 0, as no injuries are anticipated in the event of failure. Financial impact carries a weight of 0.25 and a criticality of 0.25, signifying localised equipment damage. The informational dimension, also weighted at 0.25, presents a higher criticality of 0.75 owing to the component’s role in managing system configurations that could substantially affect overall CPS security. Operationally, the component is heavily weighted (0.75) with a matching criticality of 0.75, as it directly influences CI operations – its failure potentially leading to cascading outages or grid instability. Finally, the staging impact holds a moderate weight of 0.5 and a criticality of 0.5, denoting the component’s potential to facilitate attacker persistence or lateral movement.
@@
Finally, a multi-optimisation study is performed to identify Pareto-optimal countermeasure portfolios. The scatter plot (Figure 3) reveals a consistent curved slope across the likelihood and impact axes, with data points forming a relatively smooth trajectory that appears to ascend with increasing impact levels. This suggests a stable correlation between these dimensions, wherein higher-impact and higher-availability scenarios tend to co-occur. The continuity of the slope implies underlying systemic regularities, potentially driven by latent causal BN-represented structures within the CI environment. The scatter plot also displays a plateau between approximately 0.25 and 0.30 along the Impact axis. This observation indicates that a subset of countermeasure portfolios successfully reduce the likelihood of severe impact, yet offer limited gains in system availability.
@@ 3.6 Frequency-Based Heuristic Analysis
Lastly, a frequency-based heuristic analysis of Pareto-optimal portfolios is conducted across multiple optimisation runs to rank and identify the vulnerability nodes that most consistently and effectively contributed to high-performing countermeasure configurations. These nodes are strong candidates for prioritisation in mitigation strategies, particularly under time and resource constraints during incident response.
@@
As shown in Figure 4, vulnerabilities V1 through V4 demonstrate consistently high mitigation effectiveness across numerous portfolio instances. This finding is corroborated by the underlying Bayesian Network structure (Figure 2), which reveals that these vulnerabilities constitute critical entry points, facilitating downstream attack propagation throughout the CI.
@@
  • Optimisation results reveal systemic regularity, suggesting latent causal structures within the BN structure.
@@
  • A subset of countermeasure portfolios successfully reduce the likelihood of severe impact, yet offer limited gains in system availability.
@@
  • Vulnerabilities V1V4 demonstrated consistently high mitigation effectiveness, reflecting their positions as critical entry points in the BN structure.
@@
Taken together, these findings point to several practical implications for CI practitioners. The observed systemic regularity invites further investigation into causal structures, potentially enabling more targeted interventions. Furthermore, the limited availability gains observed suggest the need for integrating complementary strategies such as anomaly detection, automated containment, or real-time system reconfiguration to strengthen operational resilience. The consistent effectiveness of vulnerabilities V1V4 reinforces their prioritisation as part of any initial mitigation sweep, especially in scenarios involving time- and resource-constrained decision-making.
4 Further Evaluation Across CI Scenarios
@@
This section evaluates the practical utility and applicability of our proposed framework by applying the methodology to another two representative CI scenarios beyond the initial solar PV inverter case study. The corresponding results are presented and analysed to highlight key insights, limitations, and implications for real-world deployment.
4.1 Ukrainian Power Grid (BlackEnergy) Attack
@@
Our proposed framework is applied to enable real-time, adaptive mitigation of the 2015 Ukrainian power grid attack, which notoriously leveraged the BlackEnergy malware to attack a CPS [36]. Figure 5 presents the corresponding Bayesian Network graph of the attack tree [27] created with the GeNIe Academic tool [12]. Based on this structure, the BN model is constructed using available information to define the asset, vulnerability, and hazard nodes, and encoded in AutomationML [4].
Figure 5: Bayesian Network of the Ukrainian power grid (BlackEnergy) attack, adapted from the attack tree model by Kumar et al. [27]. Node colours denote assets (orange), vulnerabilities (green), hazards (yellow), attack feasibility (cyan), and attacker goal (red). The graph shows how exploits and attack procedures propagate through CPS assets to cause grid disruption and support probabilistic risk analysis. A high-resolution version is available in the project’s GitHub repository ([20]).
Next, we calculate the exposure and impact probability values using the methodology outlined in Section 3. Two of the vulnerability nodes are listed in the National Vulnerability Database (NVD) [32] with CVE identifiers; the rest are not. Table 4 presents the vulnerability metrics for vulnerability node V2 (CVE-2014-1761: Microsoft Word Remote Code Execution). Equation 2 yields a final maximum exposure probability of $P(E)_{V2,CVE}=1.0$.
Table 4: Vulnerability metrics for CVE-2014-1761 (retrieved 2026-02-15). Despite age (>>10 years), high CVSS (7.8), LEV (1.0), and KEV listing yield maximum exposure probability (1.0) – indicating the vulnerability is almost certain to be exploited in real-world scenarios.
Scoring System Score
CVSS v3.1 7.8
EPSS (Latest Score) 0.9313
LEV 1.000 (100%)
KEV 1 (Listed in CISA catalogue)
$P(E)_{\text{V3, CVE}}$ 1.0
@@
Next, the exposure probabilities of non-CVE vulnerabilities must be estimated and calibrated. Taking the example of one such vulnerability node V3, which represents a malware downloader, the CVSS-proxy vector is approximated as:
@@
This CVSS vector characterises a local attack scenario within the critical infrastructure: AV:L indicates the attacker requires local access to the target asset; AC:L denotes low complexity with no specialised preconditions; PR:N and UI:N indicate that exploitation requires no privileges or user interaction; S:U shows the impact is limited to the same system; the C:H, I:H, and A:H designations reflect high levels of confidentiality, integrity, and availability loss, respectively, implying complete system compromise, including access to sensitive data, arbitrary control, and potential operational failure. Based on the CVSS vector, the fallback exposure probability estimate is derived as:
$P(E)_{V3,non-CVE}=0.306$ (19)
@@
Table 5 presents the conditional probabilities for exposure and impact associated with the attacker’s goal: H5_Disable_Electrical_Supply, derived using Bayesian inference via variable elimination. The results show a posterior likelihood of successful attack of $P(E)=0.8750$ (using an AF modifier of $\phi=1.0$), and a corresponding posterior probability of severe impact on the CPS of $P(I)=0.2393$. The composite risk score is computed as $R=P(E)\times P(I)=20.94\%$ in the BlackEnergy attack.
@@
Next, we performed optimisation studies to search for Pareto-optimal countermeasure portfolios to support decision-making in mitigating the BlackEnergy attack. Figure 6 reveals an increasingly linear relationship between impact probability and system availability at higher impact probability levels. Additionally, the Pareto fronts appear clustered near a similar value along the attack likelihood axis, suggesting limited variation in attack likelihood across trials.
Analysis of the top-performing countermeasure portfolio presented in Table 6 highlights a clear trade-off between exposure, impact, and system availability. The model prescribes near maximal intervention across all identified vulnerabilities, with mitigation parameters approaching unity. Despite these intensive efforts, the probability of severe impact remains moderate. This suggests that in the context of the BlackEnergy attack, aggressive mitigation successfully bolsters system availability but fails to fully suppress the residual risk of severe impact on the CPS. Such a result indicates that structural interdependencies within CIs may sustain severe impact even when individual node exposure is significantly reduced.
@@
In summary, the key findings from the framework’s application to the Ukrainian power grid attack are:
  • Increasingly linear relationship between impact probability and system availability at higher impact probability levels.
@@
  • Aggressive mitigation improves availability, but does not fully suppress the residual risk of severe impact on the CI.
@@
These insights equip CI practitioners with the rationale to adopt a multi-pronged mitigation strategy tailored to the dynamics of BlackEnergy-like threats. Firstly, CI practitioners should invest in early-stage interventions to reduce exposure before the system enters high-risk zones. Secondly, the clustering of Pareto fronts around similar attack likelihood values calls for prioritising interventions based on impact severity and system availability, rather than likelihood alone. Even when near-total mitigation coverage is pursued aggressively, residual risks may persist. This warrants the need to complement countermeasure portfolios with real-time anomaly detection, automated containment mechanisms, or fail-safe protocols to arrest cascading failures. Finally, CI practitioners should adopt mitigation strategies designed to target multiple vulnerabilities in parallel to enhance CI resilience.